# Parameter-Efficient Fine-Tuning via Circular Convolution

### v2

@@ 4.2 Natural Language Understanding
Results. Results are summarized in Table 2. In general, C3Ab=768/1 and C3Ab=1024/1 achieves better or on-par performance compared to baseline methods with extremely small trainable parameters. When a larger number of trainable parameters is allowed, C3Ab=768/6 and C3Ab=1024/8 significantly outperform baselines. In addition, compared to (IA)3, LoRA, VeRA, and BOFT, despite training a similar number of or more parameters, C3A stands out for its excellent memory efficiency. The only method with better memory efficiency is BitFit, which can be seen as an upper bound since it does not introduce any new parameters. Furthermore, nearly all circulant delta matrices discovered by C3A are full rank, indicating maximal capacity Zeng & Lee (2023) and providing theoretical support for the impressive results.
4.3 Instruction Tuning
Settings. We evaluate C3A against LoRA by fine-tuning LLaMA2-7B/13B and LLaMA3-8B on 4 standard instruction tuning datasets: GSM8k Cobbe et al. (2021), MATH Yu et al. (2023), ViGGO Juraska et al. (2019), and SQL Zhong et al. (2017), following the setting outlined in Nikdan et al. (2024). For the SQL dataset, we preprocess by selecting 12.5% of the dataset, followed by an 4:1 train-test split. To ensure a fair comparison, we maintain LoRA parameters with $r=32$ and a dropout rate of 0.05, while exploring various $\alpha$ values ranging from 32 to 128. All experiments are conducted for a standard single pass (epoch) over the dataset.
Results. In Table 3, we summarize our main experiments. It can be observed that C3A generally surpasses LoRA across both LLaMA2-7B/13B and LLaMA3-8B. Notably, C3A only utilizes less than half of the parameter count compared to LoRA. Overall, these results clearly highlight the effectiveness of C3A.
Table 3: Comparison of C3A and LoRA on fine-tuning LLaMA2 and LLaMA3 models in terms of accuracy and trainable parameters. The best results for each dataset are highlighted in bold. “Avg.” denotes the average accuracy of each method across all datasets.
Model Method # Trainable Parameters GSM8k MATH ViGGO SQL Avg.
LLaMA2-7B LoRAr=32 16.8M 27.22 4.10 92.24 65.39 47.24
C3Ab=4096/32 8.4M 29.34 4.25 94.03 66.21 48.46
LLaMA2-13B LoRAr=32 26.2M 42.68 6.70 93.19 67.53 52.53
C3Ab=5120/32 13.1M 42.00 6.65 95.57 67.28 52.88
LLaMA3-8B LoRAr=32 13.6M 56.57 18.95 91.78 68.19 58.87
C3Ab=4096/32 5.2M 57.54 19.75 93.07 69.47 59.96
4.4 Image Classification
@@
Settings. We focus on the task of image classification utilizing the Vision Transformer (ViT) models. Specifically, we employ both the Base and Large variants of this prevalent computer vision foundation model, as introduced by Dosovitskiy et al. (2020). These ViT models are pre-trained on the extensive ImageNet-21K dataset Ridnik et al. (2021). For the fine-tuning phase, we utilize a diverse set of datasets including Pets, Cars, DTD, EuroSAT, FGVC and RESISC.
Table 4: Fine-tuning results with ViT-Base and ViT-Large models on various image classification datasets. The models are fine-tuned for 10 epochs, and the best-performing model, based on validation set accuracy, is selected. The reported accuracy corresponds to the performance on the test set. The best results for each dataset are highlighted in bold. “Avg.” denotes the average accuracy of each method across all datasets.
Method # Trainable Parameters Pets Cars DTD EuroSAT FGVC RESISC Avg.
@@
LoRAr=16 1.57M 94.62$\pm$0.47 86.11$\pm$0.42 80.09$\pm$0.42 98.99$\pm$0.03 63.64$\pm$0.83 95.94$\pm$0.21 86.56
C3Ab=1024/16 786K 94.48$\pm$0.30 84.94$\pm$0.39 82.62$\pm$0.52 98.75$\pm$0.19 63.80$\pm$0.37 95.52$\pm$0.16 86.69
Results. Table 4 presents a summary of the results obtained from 6 image classification datasets using the ViT Base and Large models. The LoRA and C3A methods both show substantial performance improvements over Head Tuning, highlighting their effectiveness in the image classification domain. Notably, our method achieves comparable performance to LoRA while utilizing only half of the parameter count.
5 Conclusion
In this paper, we introduce C3A, an innovative approach to Parameter-Efficient Fine-Tuning (PEFT). Unlike LoRA, which utilizes low-rank decomposition, C3A employs circular convolution and its associated circulant matrix to model the delta weight matrix. This approach aims to decouple the rank of the delta weight matrix from the number of learnable parameters, allowing for high-rank adaptation while maintaining a limited parameter size. By leveraging the Fast Fourier Transform (FFT) in both the forward and backward passes, C3A achieves significant computational and memory efficiency. Consequently, C3A presents itself as a compelling alternative to LoRA for model fine-tuning.

### v3

@@ Results.
Results are presented in Table 2. Overall, C3Ab=768/1 and C3Ab=1024/1 achieve superior or comparable performance to baseline methods, despite using an exceptionally small number of trainable parameters. As the number of trainable parameters increases, models like C3Ab=768/6 and C3Ab=1024/8 significantly outperform the baselines. Moreover, compared to (IA)3, LoRA, VeRA, and BOFT, C3A distinguishes itself with remarkable memory efficiency. The only method with better memory efficiency is BitFit, which serves as an upper bound since it does not introduce new parameters. Furthermore, most of the delta matrices identified by C3A are of full rank, indicating maximum capacity (Zeng and Lee, 2023) and providing a theoretical basis for outstanding performance.
4.2 Instruction Fine-tuning
Commonsense Reasoning.
Our training utilizes Commonsense170K (Hu et al., 2023), aggregating multiple-choice questions from BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC-e, ARC-c (Clark et al., 2018), and OBQA (Mihaylov et al., 2018). Evaluation follows Hu et al. (2023) and Liu et al. (2024b), using greedy search and first-keyword appearance for answer determination.
Mathematical Reasoning.
We use MetaMathQA (Yu et al., 2023), containing 395K QA pairs from GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2020). Following Yu et al. (2023), we evaluate using greedy search on test sets, requiring chain-of-thought reasoning (Wei et al., 2022).
Code Generation.
Training uses Magicoder-Evol-Instruct-110k (Wei et al., 2024), a decontaminated subset of WizardCoder (Luo et al., 2024). Evaluation on HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), and their Plus variants follows EvalPlus (Liu et al., 2024a), reporting Pass@1 metrics.
Results.
In Table 3 and Table 4, our principal experimental observations are summarized. The C3A framework consistently surpasses LoRA and other baselines within the LLaMA series, with particular efficacy demonstrated in the most recent model, LLaMA3-8B. Noteworthy is the significant enhancement in the efficacy of LLaMA3-8B as a foundational model following the implementation of more sophisticated post-training techniques. This underscores the criticality of optimizing the fine-tuning protocols for this advanced model. It is also remarkable that C3A achieves such results while employing less than half the parameter count of LoRA. Furthermore, C3A achieves high rank adaptation and delivers optimal performance while maintaining remarkably low memory consumption. In contrast, VeRA and DoRA require substantially greater memory resources, and BoFT results in an OOM condition for a single H800, thereby underscoring the practicality and efficiency of C3A. Taken together, the findings robustly underscore the superior efficacy of the C3A methodology. We refer readers to Appendix D for examples of model answers after different tuning methods.
4.3 Ablation Study
Initialization.
We investigated initialization effects in C3A compared to LoRA, which is known for initialization sensitivity due to its $\mathbf{A}$ and $\mathbf{B}$ matrices (Hayou et al., 2024a). We evaluated multiple initialization strategies across five distinct tasks. Each initialization method was tested with 5 independent runs per task, producing a total of 20 runs per task. As shown in Figure 3, testing zero, Gaussian, Kaiming uniform, and Xavier uniform initializations for C3A’s convolution kernels revealed performance variations mostly within standard deviations, demonstrating C3A’s robustness to initialization choices.
Figure 3: Violin plots for runs on different tasks. Within each violin plot, a white bar indicates the range of average performance for each initialization strategy. It is clear that the choice of initialization does not affect the final result more than the intrinsic stochasticity.
Expressiveness.
We demonstrate C3A’s expressiveness on a synthetic dataset by placing 8 cluster centers on a 2D plane and sampling 30 points from their Gaussian distributions. A 3-layer MLP is used to classify these clusters. To compare the expressiveness of LoRA and C3A, we replace the middle layer with either a low-rank layer or a circulant layer, ensuring that both layers have the same number of trainable parameters for a fair comparison.
The results are presented in Figure 4. We observe that LoRAr=1 struggles with this simple classification task. In contrast, C3Ab=128/2, despite using the same number of parameters, achieves a perfect classification, comparable to a standard linear layer. This demonstrates the high expressiveness of C3A given the same parameter budget.
Figure 4: Training curves on the synthetic dataset. We refer our readers to Figure A1 in Appendix E for a visualization of the synthetic dataset.
Scaling.
To investigate the scaling effect of C3A, we examine it from both data and model perspectives, following the setup in He et al. (2025). From a data perspective, we focus on the MATH dataset and vary the training data used. As shown in Figure 5, we observe that C3A’s performance improves faster than LoRA as more data is added, suggesting it is more effective at leveraging additional data. From the model perspective, we evaluate C3A in small models (LLaMA3-8B, Mistral-7B) and large models (LLaMA3-70B, Mistral-8x7B). Results show that C3A outperforms the LoRA baseline in both settings, indicating that its benefits generalize to different model scales. Experiments on various data quantities and model sizes showcase C3A’s scalability and robustness, making it a promising approach for adapting language models.
Figure 5: Data and model scaling of C3A compared to LoRA. LM specifies LLaMA3 and MT specifies Mistral.
5 Conclusion
This manuscript introduces C3A, a novel PEFT method. Unlike LoRA’s low-rank decomposition, C3A utilizes circular convolution and circulant matrices to construct the delta weight matrix. This approach allows independent control over the delta weight matrix’s rank and the number of trainable parameters, enabling high-rank adaptation with limited parameter size. By employing FFT in forward and backward propagation, C3A achieves significant computational and memory efficiency, presenting a compelling alternative to LoRA for model fine-tuning.
6 Limitations
Despite C3A’s efficiency and effectiveness within limited parameter budgets, it has some intrinsic drawbacks. First, due to the requirement of $b\in\mathrm{gcd}(d_{1},d_{2})$, C3A may not be flexible enough to tackle all transformer architectures. Besides, the integration of FFT in C3A necessitates specialized kernel support for optimal performance. While this support is well-developed and readily available for NVIDIA GPUs, it may not be as mature or optimized for edge devices with different hardware architectures. This can potentially hinder the deployment and efficiency of C3A on resource-constrained edge platforms.
Appendix
Appendix A Implementations
Algorithm A1 Block-Circular Convolution PyTorch Implementation
import torch
from torch.autograd import Function
from torch.fft import fft, ifft
class BlockCircularConvolution(Function):
@staticmethod
def forward(ctx, x, w):
m, n, b = w.shape
x = x.reshape(*x.shape[:-1], n, b)
ctx.save_for_backward(x, w)
x = torch.einsum(
"...nb,mnb->...mb", ifft(x), fft(w)
x = fft(x).real
x = x.reshape(*x.shape[:-2], -1)
return x
@staticmethod
def backward(ctx, grad_output):
x, w = ctx.saved_tensors
m, n, b = w.shape
grad_output = grad_output.reshape(
*grad_output.shape[:-1], m, b
grad_output_fft = fft(grad_output)
x_grad = fft(torch.einsum(
"...mb,mnb->...nb",
grad_output_fft, ifft(w)
)).real
x_grad = x_grad.reshape(
*x_grad.shape[:-2], -1
w_grad = fft(torch.einsum(
"...mb,...nb->mnb",
grad_output_fft, ifft(x)
)).real
return x_grad, w_grad
We present the PyTorch implementation of Block-Circular Convolution in Algorithm A1. Furthermore, due to the inefficiency of directly assigning entries (as shown in Equation 4), we derive an alternative algorithm to compute the $\Delta\mathbf{W}$ more efficiently. Rather than direct assignment, we employ a forward process on the Identity matrix. Mathematically, this can be expressed as
$\Delta\mathbf{W}$ $=\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})$
$=\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\cdot\mathbf{I}_{d_{2}}$
$=\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\cdot\left[\mathbf{e}_{1},\mathbf{e}_{2},\cdots,\mathbf{e}_{d_{2}}\right]$
$=\left[\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\mathbf{e}_{1},\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\mathbf{e}_{2},\cdots,\mathcal{C}_{\mathrm{blk}}(\Delta\mathbf{w})\mathbf{e}_{d_{2}}\right]$
$=\left[\Delta\mathbf{w}\star\mathbf{e}_{1},\Delta\mathbf{w}\star\mathbf{e}_{2},\cdots,\Delta\mathbf{w}\star\mathbf{e}_{d_{2}}\right].$
Where $\mathbf{I}_{d_{2}}\in\mathbb{R}^{d_{2}\times d_{2}}$ represents an Identity matrix and $\mathbf{e}_{i}$ is the $i$th column of it. In pytorch, we can efficiently compute the iFFT of $\{\mathbf{e}_{i}\}_{i=1,2,\cdots,d_{2}}$ by a column-wise iFFT of $\mathbf{I}_{d_{2}}$. We present the Pytorch implementation in Algorithm A2 as well.
Algorithm A2 Fast Algorithm of Getting $\Delta\mathbf{W}$
import torch
from torch.fft import fft, ifft
def get_circulant_fast(w):
m, n, b = w.shape
x = torch.eye(n*b)
x = x.reshape(*x.shape[:-1], n, b)
x = torch.einsum(
"...nb,mnb->...mb", ifft(x), fft(w)
x = fft(x).real.flatten(start_dim=1).T
return x
Appendix B Image Classification
Settings.
@@ Appendix B Image Classification
In this study, we concentrate on the task of image classification leveraging Vision Transformer (ViT) models. Specifically, we employ both the Base and Large variants of this prominent foundational computer vision model, as delineated by (Dosovitskiy et al., 2020). These ViT models undergo pre-training on the expansive ImageNet-21K dataset (Ridnik et al., 2021). During the fine-tuning phase, we use an eclectic array of datasets encompassing Pets (Parkhi et al., 2012), Cars (Krause et al., 2013), DTD (Cimpoi et al., 2014), EuroSAT (Helber et al., 2019), FGVC (Maji et al., 2013), and RESISC (Cheng et al., 2017). Statistics for these datasets are provided in Table A1.
Dataset #Train #Validation #Test #Class
Pets (Parkhi et al., 2012) 3,312 368 3,669 37
Cars (Krause et al., 2013) 7,329 815 8,041 196
DTD (Cimpoi et al., 2014) 4,060 452 1,128 47
EuroSAT (Helber et al., 2019) 16,200 5,400 5,400 10
FGVC (Maji et al., 2013) 3,000 334 3,333 100
RESISC (Cheng et al., 2017) 18,900 6,300 6,300 45
Table A1: Details about the vision datasets.
Method # Params Pets Cars DTD EuroSAT FGVC RESISC Avg.
@@
LoRAr=16 1.57M 94.62$\pm$0.47 86.11$\pm$0.42 80.09$\pm$0.42 98.99$\pm$0.03 63.64$\pm$0.83 95.52$\pm$0.21 86.56
C3Ab=1024/16 0.79M 94.48$\pm$0.30 84.94$\pm$0.39 82.62$\pm$0.52 98.75$\pm$0.19 63.80$\pm$0.37 95.94$\pm$0.16 86.69
Table A2: Fine-tuning results with ViT-Base and ViT-Large models on various image classification datasets. The models are fine-tuned for 10 epochs, and the best-performing model, based on validation set accuracy, is selected. The reported accuracy corresponds to the performance on the test set. The best results between LoRA and C3A for each dataset are highlighted in bold. “Avg.” denotes the average accuracy of each method across all datasets.
Results.
Table A2 delineates a comprehensive summary of the outcomes derived from six distinct image classification datasets employing the ViT Base and Large models. The LoRA and C3A techniques exhibit significant enhancements in performance relative to Head Tuning, thereby underscoring their efficacy within the realm of image classification. Remarkably, our methodology demonstrates a performance on par with LoRA while necessitating only half of the parameter count.
Appendix C GLUE Benchmark Details
Corpus Task # Train # Val # Test # Labels Metrics Domain
Single-Sentence Tasks
CoLA Acceptability 8.55k 1.04k 1.06k 2 Matthews Corr. misc.
SST-2 Sentiment 67.3k 872 1.82k 2 Accuracy Movie reviews
Similarity and Paraphrase Tasks
MRPC Paraphrase 3.67 408 1.73k 2 Accuracy/F1 News
STS-B Sentence similarity 5.75k 1.5k 1.38k 1 Pearson/Spearman Corr. misc.
QQP Paraphrase 364k 40.4k 391k 2 Accuracy/F1 Social QA
Inference Tasks
MNLI NLI 393k 19.65k 19.65k 3 Accuracy misc.
QNLI QA/NLI 105k 5.46k 5.46k 2 Accuracy Wikipedia
RTE NLI 2.49k 277 3k 2 Accuracy News & Wikipedia
Table A3: Task descriptions and dataset statistics of the GLUE benchmark (Wang et al., 2018).
The General Language Understanding Evaluation (GLUE) benchmark is a collection of nine natural language understanding tasks designed to test a model’s performance on a diverse set of natural language understanding challenges. It was introduced by researchers from New York University, the University of Washington, and DeepMind in a 2018 paper titled “GLUE: A Multi-Task Benchmark and Analysis Platform for Natural Language Understanding” (Wang et al., 2018).
The nine tasks included in the benchmark are:
  1. 1.
    CoLA (Corpus of Linguistic Acceptability): A binary classification task that tests a model’s ability to determine whether a given sentence is grammatically acceptable.
  2. 2.
    SST-2 (Stanford Sentiment Treebank): A binary sentiment classification task using movie review excerpts.
  3. 3.
    MRPC (Microsoft Research Paraphrase Corpus): A binary classification task to determine whether two sentences are semantically equivalent.
  4. 4.
    STS-B (Semantic Textual Similarity Benchmark): A regression task that scores the semantic similarity of sentence pairs on a scale from 1 to 5.
  5. 5.
    QQP (Quora Question Pairs): A binary classification task to determine whether two questions are semantically equivalent.
  6. 6.
    MNLI (Multi-Genre Natural Language Inference): A ternary classification task that tests a model’s ability to perform textual entailment across multiple genres.
  7. 7.
    QNLI (Question Natural Language Inference): A binary classification task that tests a model’s ability to determine whether a sentence contains the answer to a given question.
  8. 8.
    RTE (Recognizing Textual Entailment): A binary classification task that tests a model’s ability to determine whether a premise entails a hypothesis.
  9. 9.
    WNLI (Winograd Natural Language Inference): A binary classification task that tests a model’s ability to resolve ambiguous pronouns in a sentence.
The GLUE benchmark has become a widely-used standard for evaluating the performance of pre-trained language models and fine-tuned models on a variety of natural language understanding tasks. The benchmark provides a single-number metric, the GLUE score, which is an average of the scores on each individual task. This allows for easy comparison of different models and architectures. We refer readers to Table A3 for more details.
Appendix D Instruction Fine-tuning Examples
Example D.1: MATH Below is an instruction that describes a task. Write a response that appropriately completes the request. Instruction: Find the number of square units bounded by the $x$-axis, the $y$-axis and the line $x+y=6$. Answer the above question. First think step by step and then answer the final number. LoRA Answer: The line $x+y=6$ intersects the $x$-axis at $x=6$ and the $y$-axis at $y=6$. So, the line divides the square with side length 6 into two triangles. The area of each triangle is $\frac{1}{2}\times 6\times 6=18$. So, the area of the square is $2\times 18=\boxed{36}$. The final answer is: 36. C3A Answer: The line $x+y=6$ intersects the $x$-axis at $(6,0)$ and the $y$-axis at $(0,6)$. Thus, the line divides the first quadrant into two triangles, one with base $6$ and height $6$ and the other with base $6$ and height 00. The area of the first triangle is $(6\times 6)/2=18$ square units, and the area of the second triangle is $(6\times 0)/2=0$ square units. Thus, the total area is $18+0=\boxed{18}$ square units. The final answer is: 18.
Example D.2: GSM8K Below is an instruction that describes a task. Write a response that appropriately completes the request. Instruction: Tim makes a special honey and jam mix. To make 1 jar of the mixture it takes 2 pounds of passion fruit and 1.5 pounds of jackfruit. It also takes 2 pounds of special honey. The passion fruit costs \$6 per pound, the jackfruit is \$8 per pound, and the honey is \$10 per pound. He sells each jar for \$50. How much profit does he make selling 10 jars? Answer the above question. First think step by step and then answer the final number. LoRA Answer: The passion fruit costs $6\times 2=12$. The jackfruit costs $8\times 1.5=12$. The honey costs $10\times 2=20$. So the total cost is $12+12+20=44$. So he makes a profit of $50\times 10-44=\boxed{460}$. The final answer is: 460. C3A Answer: The passion fruit costs $6\times 2=12$. The jackfruit costs $8\times 1.5=12$. So the fruit costs $12+12=24$. The honey costs $10\times 2=20$. So the total cost per jar is $24+20=44$. So he makes a profit of $50-44=6$ per jar. So he makes a profit of $6\times 10=\boxed{60}$. The final answer is: 60.
Appendix E Synthetic Dataset
We visualize the synthetic dataset we use for the expressiveness ablation study in Figure A1.
Figure A1: Synthetic dataset used for the expresiveness ablation study.
Appendix F Hyperparameters
Model Hyperparameter SST-2 MRPC CoLA QNLI RTE STS-B
Both Optimizer AdamW
LR Schedule Linear
Warmup Ratio 0.06
C3A Initialization Xavier Uniform
Max Seq. Len 512
Base Epochs 40 80 80 40 80 80
Batch Size 128 128 128 64 64 128
Learning Rate (C3Ab=768/6) 2E-1 3E-1 2E-1 7E-2 3E-1 2E-1
Learning Rate (Head) 2E-4 4E-6 3E-2 8E-6 6E-3 4E-2
Large Epochs 10 80 70 30 60 40
Batch Size 128 128 128 32 64 128
Learning Rate (C3Ab=1024/8) 9E-2 3E-1 2E-1 7E-2 5E-2 2E-1
Learning Rate (Head) 2E-4 5E-6 3E-3 8E-6 3E-3 5E-4
Table A4: Hyperparameter setup of C3A for the GLUE benchmark.
Model Hyperparameter Pets Cars DTD EuroSAT FGVC RESISC
Both Optimizer AdamW
LR Schedule None
C3A Initialization Xavier Uniform
Epochs 10
Batch Size 64
Base Learning Rate (C3Ab=768/12) 4E-1 4E+0 2E+0 2E+0 7E+0 2E+0
Learning Rate (Head) 1E-2 1E-2 2E-2 8E-3 1E-2 2E-2
Weight Decay 3E-4 5E-4 6E-5 2E-5 1E-5 2E-5
Large Learning Rate (C3Ab=1024/16) 7E-1 4E+0 2E+0 2E+0 4E+0 3E+0
Learning Rate (Head) 3E-3 8E-3 7E-3 2E-2 1E-1 4E-3
Weight Decay 4E-3 1E-5 2E-4 5E-4 2E-5 9E-5
Table A5: Hyperparameter setup of C3A for image classification tasks.
Hyperparameter Common Math Code
Optimizer AdamW
LR Scheduler Cosine
Batch Size 16 128 128
Warmup Ratio 0.05 0.03 0.03
Epoch 3 2 2
Dropout 0.05
Learning Rate (LLaMA2-7B) 3E-1 4E-1 6E-1
Learning Rate (LLaMA3-8B) 3E-1 5E-1 6E-1
Table A6: Hyperparameter setup of C3A for instruction tuning.